Use the following information for the next four problems.  Researchers believe that body temperatures for a particular...

  1. Use the following information for the next four problems.  Researchers believe that body temperatures for a particular population vary according to a normal model. They want to estimate , the mean body temperature for that population. They randomly select a group of 17 adults and record the body temperature for each person. That data results in a sample mean of 98.3 F and a sample standard deviation of 0.68 F.

    Calculate a 90% confidence interval for estimating the mean.

    a.

    98.3  0.29

    b.

    98.3  0.07

    c.

    98.3  0.27

    d.

    98.3  0.15

  2. Which two of the following conditions must be true in order for the confidence interval above to be valid?

    a.

    The population distribution is normal (or nearly normal) in shape.

    b.

    The sample size  is large ().

    c.

    The sample observations are independent of each other.

    d.

    and

  3. A 90% confidence interval based on a smaller sample would be __________ the confidence interval calculated above.

    a.

    the same length as

    b.

    shorter than

    c.

    longer than

  4. Suppose that the confidence interval above had ranged from 97.9 to 98.7. Which one of the following would give a correct interpretation of the interval?

    a.

    90% of the people in the sample have body temperatures between 97.9 and 98.7.

    b.

    We are 90% confident that the mean body temperature of all people in the population is between 97.9 and 98.7.

    c.

    We are 90% confident that the mean body temperature of the people in the sample is between 97.9 and 98.7.

    d.

    90% of all people in the population have body temperatures between 97.9 and 98.7.
Answer #1

x=98.3 S = 0.60 n=17 2= 1-0.90, = 0.10 !! . Find tarzin-1 = to.107, 917-1 = toos, 16 = 1.746 cas calculate 904. confidence in